OpenStudy (anonymous):

\log_{2}(x-3) + \log_{2}(x) - \log_{2}(x+2) = 2

OpenStudy (anonymous):

\[\log_{2}(x-3) + \log_{2}(x) - \log_{2}(x+2) = 2 \]

OpenStudy (anonymous):

\[\log_2((x-3)x)-\log_3(x+2)=2\] \[\log_2(\frac{x(x-2)}{x+2})=2\] \[\frac{x(x-2)}{x+2}=2^2=4\] \[x(x-2)=4(x+2)\] and now solve the quadratic

OpenStudy (anonymous):

\[x^2-6x-8=0\] \[x=3\pm\sqrt{17}\] \[x=3+\sqrt{17}\] because the other one is negative